# Deconfined quantum criticality of frustrated hard-core dipolar bosons

### v2

@@ Abstract
Deconfined quantum critical points (DQCPs) are proposed as unconventional second-order phase transitions beyond the Landau-Ginzburg-Wilson paradigm. The nature and experimental realizations of DQCPs are crucial issues of importance. We illustrate the potential for DQCPs between the valence bond solid state and the antiferromagnetic phase to arise in optical lattices containing frustrated dipolar bosons subject to hard-core constraints. The emergence of DQCPs is comprehended through the fusion of two Berezinskii-Kosterlitz-Thouless (BKT) transitions. The DQCPs and the BKTs are confirmed by the finite-size scaling of ground-state fidelity susceptibilities. The numerical analysis reveals varying critical exponents of the correlation length in DQCPs and the logarithmic scaling in BKTs, respectively. This work offers a promising platform for realizing DQCPs and provides valuable insights into their nature within the framework of topological phase transitions.
Introduction-. The quantum phase transition is one of the fundamental concepts in realms of statistical physics and condensed matter physics [1]. Continuous phase transitions are effectively characterized by the Landau-Ginzburg-Wilson (LGW) paradigm through the utilization of local order parameters based on the concept of spontaneous symmetry breaking. Berezinskii–Kosterlitz–Thouless (BKT) transitions [2, 3] and deconfined quantum critical points (DQCPs) [4, 5] stand as two renowned examples that extend beyond the framework of the LGW theory. BKT transitions are infinite-order phase transitions that occur between quasi-ordered phases and disordered phases [2, 3], while DQCPs represent second-order phase transitions between two ordered phases [4, 5]. Over last decades, DQCPs have garnered significant attention, particularly in exploring the nature of phase transitions (first or second order) [4, 5, 6], quantum criticalities [7], duality phenomena [8, 9, 10, 11], and emergent symmetries [11, 12, 13]. In previous studies, BKT transitions and DQCPs were investigated separately. Here, we demonstrate that DQCPs can arise from the fusion of two BKT transitions, offering essential insights into comprehending both BKT transitions and DQCPs.
@@
In this letter, we instead explore a frustrated dipolar Bose-Hubbard (FDBH) model in one dimension under the hard-core limit, where nearest-neighbor interactions stem from dipolar interactions [36], which are achievable within an atomic mixture [35] as well. We show that the FDBH model hosts DQCPs, aligning with recent theoretical proposals [13, 37, 38, 39, 40] that describe a continuous quantum phase transition from an antiferromagnetic (AFM) state to a valence bond solid (VBS) state. We employ the density matrix renormalization group (DMRG) technique [41, 42, 43, 44, 45] to investigate quantum phase transitions of FDBH model by varying nearest-neighbor interactions. Our study reveals the presence of three distinct phases, delineated by two BKT transitions and one DQCP through the fidelity susceptibility. Intriguingly, our proposal reveals that DQCPs emerge from the fusion of two BKT transitions, providing an essential insight into the understanding of both BKT transitions and DQCPs.
Model-. We consider interacting dipolar bosons within a one-dimensional lattice accounting for both nearest-neighbor and next-nearest-neighbor hopping terms, as illustrated in Fig.1(a). The corresponding Hamiltonian of the FDBH model is described by
@@
In this Hamiltonian, $b_{r}$ ($b_{r}^{\dagger}$) denote the bosonic annihilation (creation) operators at the $r$th site. And, $n_{r}=b_{r}^{\dagger}b_{r}$ represents the particle number operator, while $t_{1}\geq 0$ and $t_{2}\geq 0$ denote the amplitudes of the nearest-neighbor and next-nearest-neighbor hopping, respectively. Moreover, $V\geq 0$ and $U\geq 0$ signify the strengths of the nearest-neighbor and on-site interactions.
@@
The frustrated one-dimensional chain is equivalent to the zigzag chain [cf. Fig. 1(b)], which can be formed by the incoherent superposition of an optical triangular lattice [24]. Ultracold bosons offer a wealth of physics arising from the interplay between frustrations induced by lattice geometry and interactions [24]. For instance, it is demonstrated that ultracold bosons subject to a three-body constraint exhibit chiral superfluid and Haldane insulator phases within zig-zag optical lattices [24]. In the hard-core limit ($U\rightarrow\infty$), the FDBH model in Eq.(1) becomes
@@
Figure 3: Phase transition between the XY phase and the VBS phase in hard-core FDBH chain at $V/t_{1}=0.1$ in OBCs. (a) Fidelity susceptibility $\chi_{L}$ is examined with respect to $t_{2}/t_{1}$, in which the peak values and positions of $\chi_{L}$ gradually increase and shift to the left, respectively, as the system size increases. (b) The VBS order parameter $O_{\text{VBS}}$ plotted against $t_{2}/t_{1}$ for $L=192$. The red dashed line indicates the critical point $t_{2c}/t_{1}$. (c) The finite-size scaling of the peak values of $\chi_{L}^{m}$ with respect to the system sizes $1/\ln{(aL)}$ confirms the BKT transition, with $a$ representing the cutoff constant. (d) The scaling of the peak positions $t_{2,m}/t_{1}$ as a function of the system sizes $1/\ln^{2}{(aL)}$, where the critical value obtained through the extrapolation is $t_{2c}/t_{1}=0.4107$, further confirming the presence of the BKT transition.
Berezinskii-Kosterlitz-Thouless transitions-. Subject to the hard-core constraints ($U\rightarrow\infty$), the FDBH model described in Eq.(2) can be effectively transformed into a spin model, as presented below:
@@
Figure 4: Phase transition between the XY phase and the AFM phase in hard-core FDBH chain at $t_{2}/t_{1}=0.1$. (a) The fidelity susceptibility $\chi_{L}$ in OBCs varies with $V/t_{1}$, exhibiting an increase in peak values and leftward shifts in peak positions with increasing system size. (b) The AFM order parameter $O_{\text{AFM}}$ plotted against $V/t_{1}$ for $L=96$ in PBCs. The red dashed line indicates the critical point $V_{c}/t_{1}$. (c) The finite-size scaling of the peak values of $\chi_{L}^{m}$ with respect to the system sizes $1/\ln{(aL)}$ confirms the BKT transition, with $a$ representing the cutoff constant. (d) The scaling of the peak positions $V_{m}/t_{1}$ as a function of the system sizes $1/\ln^{2}{(aL)}$, where the critical value obtained through the extrapolation is $V_{c}/V_{1}=1.729$, further confirming the presence of the BKT transition.
When $V=0$, the system shown in Eq.(3) become the XY model incorporating both the nearest-neighbor and next-nearest-neighbor hopping terms. In particular, when $t_{2}=0$, the system’s ground state is characterized by the gapless XY (superfluidity (SF) in bosonic language) phase (see Supplemental Material [46] for details). Upon increasing $t_{2}$, the system undergoes a phase transition from the XY phase to the VBS (bond-ordered-wave (BOW) in bosonic language) phase [35, 47, 48, 49, 50]. To verify if the transition between the XY phase and the VBS phase persists in the presence of interaction $V$, we calculate the ground-state phase diagram of the FDBH model described in Eq.(2) varying parameters $t_{2}/t_{1}$ and $V/t_{1}$ in terms of DMRG method. During DMRG simulations, we keep 500 matrix states and enforce $n_{\text{max}}=1$ per site. The full phase diagram is shown in Fig.2, which is obtained from the fidelity susceptibility [51, 52, 53, 54, 55, 56, 57, 58],
$\chi_{L}=\frac{1}{L}\lim\limits_{\delta\lambda\to 0}\frac{-2\ln F(\lambda,\lambda+\delta\lambda)}{(\delta\lambda)^{2}},$ (4)
where $F(\lambda,\lambda+\delta\lambda)=|\langle\Psi_{0}(\lambda)|\Psi_{0}(\lambda+\delta\lambda)\rangle|$ is the fidelity with the control parameter $\lambda=\{t_{2}/t_{1},V/t_{1}\}$. The numerical results support the argument that the phase transition persists even under weak interaction. In Fig.3(a), we depict the scaling behavior of fidelity susceptibility at an interaction strength of $V/t_{1}=0.1$ by varying the next-nearest-neighbor hopping $t_{2}/t_{1}$. It is observed that the fidelity susceptibility increases gradually as the system’s size increases. Furthermore, both the maximum fidelity susceptibility $\chi_{L}^{m}$ and the peak positions $\lambda_{m}$ themselves are demonstrated to be in excellent agreement with the scaling laws [c.f. Fig.3(c) and (d)] proposed for BKT transitions [59], as described by:
@@
where $\chi_{\infty}$ represents the fidelity susceptibility in the thermodynamic limit, $\lambda_{c}$ is the critical point, and $a$ is a positive nonuniversal constant serving as a cutoff. The transition from the XY phase to the VBS phase is further validated through level crossing (see Supplemental Material [46] for details) and the order parameter [c.f. Fig.3(b)] characterizing the VBS phase under open boundary conditions, as expressed in:
@@
Figure 5: Phase transition between the VBS phase and the AFM phase in hard-core FDBH chain at $t_{2}/t_{1}=0.7$. (a) The fidelity susceptibility $\chi_{L}$ varies with $V$ in OBCs. The peak position of $\chi_{L}$ shifts to the right as the system size increases. (b) The VBS and AFM order parameters $O_{\text{VBS}}$ and $O_{\text{AFM}}$ plotted against $V/t_{1}$ for $L=192$ in OBCs and $L=96$ in PBCs, respectively. The red dashed line denotes the critical point $V_{c}/t_{1}$. (c) The finite-size scaling of the peak values of $\ln(\chi_{L}^{m})$ with respect to the system sizes $\ln(L)$ confirms the 2nd transition with the correlation length critical exponent $\nu=1.1112$. (d) The scaling of the peak positions $V_{m}/t_{1}$ as a function of the system sizes $1/L$, where the critical value obtained through the extrapolation is $V_{c}/V_{1}=2.122$, further confirming the presence of the 2nd transition.
When $t_{2}=0$, the model in Eq.(3) represents the quantum spin-1/2 antiferromagnetic XXZ chain, which undergoes a BKT transition from the XY phase ($V/t_{1}<2$) to the AFM (density wave (DW) with all particles located in one of the legs in the bosonic language) phase ($V/t_{1}>2$) [59]. Our numerical results obtained from the fidelity susceptibility confirm that the critical value is located at $V/t_{1}=2$, as illustrated in Fig.2. We then investigate whether this phase transition persists under the next-nearest-neighbor hopping using the DMRG. The fidelity susceptibility is computed as a function of $V/t_{1}$ at $t_{2}=0.1$ for large open chains. The numerical results indicate that the peak of the fidelity susceptibility gradually shifts towards the left, and its magnitude slowly increases with increasing system size, as shown in Fig.4(a). Finite-size scaling analysis, presented in Fig.4(c) and (d), demonstrates that the fidelity susceptibility continues to adhere to the scaling laws proposed for BKT transitions [59], as outlined in Eq.(5) and Eq.(6). Moreover, it is observed that the long-range spin correlations in the $z$-direction under periodic boundary conditions,
$O_{\text{AFM}}$ $=\langle\sigma_{r}^{z}\sigma_{r+L/2}^{z}\rangle,$ (8)
@@
Deconfined quantum critical point-. We have demonstrated that the BKT transitions between the XY phase and VBS state, as well as the BKT transition between the XY phase and AFM phase, persist even in the presence of small next-nearest-neighbor hopping ($t_{2}/t_{1}$) and interaction ($V/t_{1}$). Now, we turn to explore the transitions in the regime characterized by large values of $t_{2}/t_{1}$ and $V/t_{1}$. Interestingly, we find unexpectedly that the system undergoes a phase transition from the AFM phase to the VBS state as $t_{2}/t_{1}$ or $V/t_{1}$ increase further.
As an example, we compute the fidelity susceptibility and the order parameters at $t_{2}/t_{1}=0.7$ [cf. Fig.5]. It is observed that the system resides in the VBS phase for $2.120<V/t_{1}<$ 2.12, characterized by a non-zero $O_{\text{VBS}}$ and vanishing $O_{\text{AFM}}$. Conversely, in the regime $32.12<V/t_{1}<$ 3, the reverse occurs, indicating that the system is in the AFM phase [cf. Fig.5(b)]. Notably, both order parameters, $O_{\text{AFM}}$ and $O_{\text{VBS}}$, seem to smoothly vanish at this same point, suggesting an unconventional nature for this quantum phase transition. This transition is argued to be a second-order phase transition with the correlation length critical exponent $\nu=1.11$ (see Supplemental Material [46] for additional critical exponents), as demonstrated in Fig.5(a), (c), and (d) from the finite-size scaling of the fidelity susceptibility [53, 54, 55],

### v3

@@ Abstract
Deconfined quantum critical points (DQCPs) are proposed as unconventional second-order phase transitions beyond the Landau-Ginzburg-Wilson paradigm. The nature and experimental realizations of DQCPs are crucial issues of importance. We illustrate the potential for DQCPs between the valence bond solid state and the antiferromagnetic phase to arise in optical lattices containing frustrated dipolar bosons subject to hard-core constraints. The emergence of DQCPs is comprehended through the fusion of two Berezinskii-Kosterlitz-Thouless (BKT) transitions. The DQCPs and the BKTs are confirmed by the scaling of ground-state fidelity susceptibilities in finite systems and the analysis of order parameters obtained from infinite systems. The numerical analysis reveals varying critical exponents of the correlation length in DQCPs and the logarithmic scaling in BKTs, respectively. This work offers a promising platform for realizing DQCPs and provides valuable insights into their nature within the framework of topological phase transitions.
I Introduction
The quantum phase transition is one of the fundamental concepts in realms of statistical physics and condensed matter physics [1]. Continuous phase transitions are effectively characterized by the Landau-Ginzburg-Wilson (LGW) paradigm through the utilization of local order parameters based on the concept of spontaneous symmetry breaking. Berezinskii–Kosterlitz–Thouless (BKT) transitions [2, 3] and deconfined quantum critical points (DQCPs) [4, 5] stand as two renowned examples that extend beyond the framework of the LGW theory. BKT transitions are infinite-order phase transitions that occur between quasi-ordered phases and disordered phases [2, 3], while DQCPs represent second-order phase transitions between two ordered phases [4, 5]. Over last decades, DQCPs have garnered significant attention, particularly in exploring the nature of phase transitions (first or second order) [4, 5, 6], quantum criticalities [7], duality phenomena [8, 9, 10, 11], and emergent symmetries [11, 12, 13]. In previous studies, BKT transitions and DQCPs were investigated separately. Here, we demonstrate that DQCPs can arise from the fusion of two BKT transitions, offering essential insights into comprehending both BKT transitions and DQCPs.
@@ I Introduction
In this paper, we instead explore a frustrated dipolar Bose-Hubbard (FDBH) model in one dimension under the hard-core limit, where nearest-neighbor interactions stem from dipolar interactions [36], which are achievable within an atomic mixture [35] as well. We show that the FDBH model hosts DQCPs, aligning with recent theoretical proposals [13, 37, 38, 39, 40] that describe a continuous quantum phase transition from an antiferromagnetic (AFM) state to a valence bond solid (VBS) state. We employ the density matrix renormalization group (DMRG) technique [41, 42, 43, 44, 45] and the infinite time evolving block decimation (iTEBD) method [46] to investigate quantum phase transitions of FDBH model by varying nearest-neighbor interactions. Our study reveals the presence of three distinct phases, delineated by two BKT transitions and one DQCP through the fidelity susceptibility. Notably, our numerical simulations confirm that DQCPs emerge from the fusion of two BKT transitions [33, 34, 35], providing an essential insight into the understanding of both BKT transitions and DQCPs.
II Model
We consider interacting dipolar bosons within a one-dimensional lattice accounting for both nearest-neighbor and next-nearest-neighbor hopping terms, as illustrated in Fig.1(a). The corresponding Hamiltonian of the FDBH model is described by
@@ II Model
In this Hamiltonian, $b_{r}$ ($b_{r}^{\dagger}$) denote the bosonic annihilation (creation) operators at the $r$th site. The expression $n_{r}=b_{r}^{\dagger}b_{r}$ represents the particle number operator, with $t_{1}\geq 0$ and $t_{2}\geq 0$ denoting the amplitudes for nearest-neighbor and next-nearest-neighbor hopping, respectively. The positive hopping amplitudes are employed to introduce frustrations [47], which can be achieved using the Floquet protocol [48, 49, 50, 51]. $V\geq 0$ and $U\geq 0$ signify the strengths of the nearest-neighbor and on-site interactions. Note that the Hamiltonian in Eq.(1) can be realized using a synthetic dimension [52, 53], where the interspecies interactions serve as dipolar interactions [35], offering an alternative approach for realizing DQCPs.
@@
The frustrated one-dimensional chain is equivalent to the zigzag chain [cf. Fig. 1(b)], which can be formed by the incoherent superposition of an optical triangular lattice [24]. Ultracold bosons offer a wealth of physics arising from the interplay between frustrations induced by lattice geometry and interactions [24]. For instance, it is demonstrated that ultracold bosons subject to a three-body constraint exhibit chiral superfluid and Haldane insulator phases within zig-zag optical lattices [24]. We note that although long-range dipolar interactions decay as $1/|r-r^{\prime}|^{3}$, the properties of hard-core dipoles in optical lattices at half-filling can be effectively understood in terms of nearest-neighbor interactions [54, 55, 23]. Therefore, we focus solely on the nearest-neighbor interaction term.
In the hard-core limit ($U\rightarrow\infty$), the FDBH model in Eq.(1) becomes
@@
Figure 3: Phase transition between the XY phase and the VBS phase in hard-core FDBH chain at $V/t_{1}=0.1$ in OBCs. (a) Fidelity susceptibility $\chi_{L}$ is examined with respect to $t_{2}/t_{1}$, in which the peak values and positions of $\chi_{L}$ gradually increase and shift to the left, respectively, as the system size increases. (b) The VBS order parameter $O_{\text{VBS}}$ plotted against $t_{2}/t_{1}$ obtained from the iTEBD method. The red dashed line indicates the critical point $t_{2c}/t_{1}$ in the thermodynamic limit from the finite scaling of the fidelity susceptibility. (c) The finite-size scaling of the peak values of $\chi_{L}^{m}$ with respect to the system sizes $1/\ln{(aL)}$ confirms the BKT transition, with $a$ representing the cutoff constant. (d) The scaling of the peak positions $t_{2,m}/t_{1}$ as a function of the system sizes $1/\ln^{2}{(aL)}$, where the critical value obtained through the extrapolation is $t_{2c}/t_{1}=0.4107$, further confirming the presence of the BKT transition.
III Berezinskii-Kosterlitz-Thouless transitions
Subject to the hard-core constraints ($U\rightarrow\infty$), the FDBH model described in Eq.(2) can be effectively transformed into a spin model, as presented below:
@@ III Berezinskii-Kosterlitz-Thouless transitions
Figure 4: Phase transition between the XY phase and the AFM phase in hard-core FDBH chain at $t_{2}/t_{1}=0.1$. (a) The fidelity susceptibility $\chi_{L}$ in OBCs varies with $V/t_{1}$, exhibiting an increase in peak values and leftward shifts in peak positions with increasing system size. (b) The AFM order parameter $O_{\text{AFM}}$ plotted against $V/t_{1}$ obtained from the iTEBD method. The red dashed line indicates the critical point $V_{c}/t_{1}$ in the thermodynamic limit from the finite scaling of the fidelity susceptibility. (c) The finite-size scaling of the peak values of $\chi_{L}^{m}$ with respect to the system sizes $1/\ln{(aL)}$ confirms the BKT transition, with $a$ representing the cutoff constant. (d) The scaling of the peak positions $V_{m}/t_{1}$ as a function of the system sizes $1/\ln^{2}{(aL)}$, where the critical value obtained through the extrapolation is $V_{c}/V_{1}=1.729$, further confirming the presence of the BKT transition.
When $V=0$, the system shown in Eq.(3) become the XY model incorporating both the nearest-neighbor and next-nearest-neighbor hopping terms. In particular, when $t_{2}=0$, the system’s ground state is characterized by the gapless XY (superfluidity (SF) in bosonic language) phase (see Appendix A for details). Upon increasing $t_{2}$, the system undergoes a phase transition from the XY phase to the VBS (bond-ordered-wave (BOW) in bosonic language) phase [35, 56, 57, 58, 59]. To verify if the transition between the XY phase and the VBS phase persists in the presence of interaction $V$, we calculate the ground-state phase diagram of the FDBH model described in Eq.(2) varying parameters $t_{2}/t_{1}$ and $V/t_{1}$ in terms of DMRG method. During DMRG simulations, we keep 500 states and enforce $n_{\text{max}}=1$ per site. The full phase diagram is shown in Fig.2, which is obtained from the fidelity susceptibility [60, 61, 62, 63, 64, 65, 66, 67],
$\chi_{L}=\frac{1}{L}\lim\limits_{\delta\lambda_{l}\to 0}\frac{-2\ln F(\lambda_{l},\lambda_{l}+\delta\lambda_{l})}{(\delta\lambda_{l})^{2}}.$ (4)
Here, $F(\lambda_{l},\lambda_{l}+\delta\lambda_{l})=|\langle\Psi_{0}(\lambda_{l})|\Psi_{0}(\lambda_{l}+\delta\lambda_{l})\rangle|$ is the fidelity for $l=1,2$, with the control parameters defined as $\lambda_{1}=t_{2}/t_{1}$ and $\lambda_{2}=V/t_{1}$. The numerical results support the argument that the phase transition persists even under weak interaction. In Fig.3(a), we depict the scaling behavior of fidelity susceptibility at an interaction strength of $V/t_{1}=0.1$ by varying the next-nearest-neighbor hopping $t_{2}/t_{1}$. It is observed that the fidelity susceptibility increases gradually as the system’s size increases. Furthermore, both the maximum fidelity susceptibility $\chi_{L}^{m}$ and the peak positions $\lambda_{m}$ themselves are demonstrated to be in excellent agreement with the scaling laws [c.f. Fig.3(c) and (d)] proposed for BKT transitions [68], as described by:
@@
where $\chi_{\infty}$ represents the fidelity susceptibility in the thermodynamic limit, $\lambda_{c}$ is the critical point, and $a$ is a positive nonuniversal constant serving as a cutoff. The transition from the XY phase to the VBS phase is further validated through level crossing (see Appendix B for details) and the order parameter [c.f. Fig.3(b)] characterizing the VBS phase under open boundary conditions (OBCs), as expressed in:
@@
Figure 5: Phase transition between the VBS phase and the AFM phase in hard-core FDBH chain at $t_{2}/t_{1}=0.7$. (a) The fidelity susceptibility $\chi_{L}$ varies with $V$ in OBCs. The peak position of $\chi_{L}$ shifts to the right as the system size increases. (b) The VBS and AFM order parameters $O_{\text{VBS}}$ and $O_{\text{AFM}}$ plotted against $V/t_{1}$ obtained from the iTEBD method. The red dashed line indicates the critical point $V_{c}/t_{1}$ in the thermodynamic limit from the finite scaling of the fidelity susceptibility. (c) The finite-size scaling of the peak values of $\ln(\chi_{L}^{m})$ with respect to the system sizes $\ln(L)$ confirms the 2nd transition with the correlation length critical exponent $\nu=1.1112$. (d) The scaling of the peak positions $V_{m}/t_{1}$ as a function of the system sizes $1/L$, where the critical value obtained through the extrapolation is $V_{c}/V_{1}=2.122$, further confirming the presence of the 2nd transition. (e) Correlation lengths as a function of $V/t_{1}$ for different numbers of states $M=64,100,200,300,400$, respectively. (f) Scaling of the entanglement entropy with respect to the correlation lengths at the critical point, yielding a central charge $c=1$.
When $t_{2}=0$, the model in Eq.(3) represents the quantum spin-1/2 antiferromagnetic XXZ chain, which undergoes a BKT transition from the XY phase ($V/t_{1}<2$) to the AFM (density wave (DW) with all particles located in one of the legs in the bosonic language) phase ($V/t_{1}>2$) [68]. Our numerical results obtained from the fidelity susceptibility confirm that the critical value is located at $V/t_{1}=2$, as illustrated in Fig.2. We then investigate whether this phase transition persists under the next-nearest-neighbor hopping using the DMRG. The fidelity susceptibility is computed as a function of $V/t_{1}$ at $t_{2}=0.1$ for large open chains. The numerical results indicate that the peak of the fidelity susceptibility gradually shifts towards the left, and its magnitude slowly increases with increasing system size, as shown in Fig.4(a). Finite-size scaling analysis, presented in Fig.4(c) and (d), demonstrates that the fidelity susceptibility continues to adhere to the scaling laws proposed for BKT transitions [68], as outlined in Eq.(5) and Eq.(6). Moreover, it is observed the staggered magnetization in the $z$-direction,
$O_{\text{AFM}}$ $=\frac{1}{L}\sum_{r}(-1)^{r}\langle\sigma_{r}^{z}\rangle,$ (8)
@@
IV Deconfined quantum critical point
We have demonstrated that the BKT transitions between the XY phase and VBS state, as well as the BKT transition between the XY phase and AFM phase, persist even in the presence of small next-nearest-neighbor hopping ($t_{2}/t_{1}$) and interaction ($V/t_{1}$). Now, we turn to explore the transitions in the regime characterized by large values of $t_{2}/t_{1}$ and $V/t_{1}$. Interestingly, we find unexpectedly that the system undergoes a phase transition from the AFM phase to the VBS state as $t_{2}/t_{1}$ or $V/t_{1}$ increase further.
As an example, we compute the fidelity susceptibility and the order parameters at $t_{2}/t_{1}=0.7$ [cf. Fig.5]. It is observed that the system resides in the VBS phase for $2.120<V/t_{1}<$ 2.12, characterized by a non-zero $O_{\text{VBS}}$ and vanishing $O_{\text{AFM}}$. Conversely, in the regime $32.12<V/t_{1}<$ 3, the reverse occurs, indicating that the system is in the AFM phase [cf. Fig.5(b)]. Notably, both order parameters, $O_{\text{AFM}}$ and $O_{\text{VBS}}$, seem to smoothly vanish at this same point, suggesting an unconventional nature for this quantum phase transition. This transition is argued to be a second-order phase transition with the correlation length critical exponent $\nu=1.11$ (see Appendix C for additional critical exponents), as demonstrated in Fig.5(a), (c), and (d) from the finite-size scaling of the fidelity susceptibility [62, 63, 64],